# Look and Talk: Seamless AI Assistant Interaction with Gaze-Triggered Activation

### v1

@@ Abstract.
Engaging with AI assistants to gather essential information in time is becoming increasingly common. Traditional activation methods, like wake words such as Hey Siri, Ok Google, and Hey Alexa, are constrained by technical challenges such as false activations, recognition errors, and discomfort in public settings. Similarly, activating AI systems via physical buttons imposes strict interactive limitations as it demands particular physical actions, which hinders fluid and spontaneous communication with AI. Our approach employs eye-tracking technology within AR glasses to discern a user’s intention to engage with the AI assistant. By sustaining eye contact on a virtual AI avatar for a specific time, users can initiate an interaction silently and without using their hands. Preliminary user feedback suggests that this technique is relatively intuitive, natural, and less obtrusive, highlighting its potential for integrating AI assistants fluidly into everyday interactions.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Interaction techniquesccs: Computing methodologies Natural language processingccs: Applied computing
@@ 1. Introduction
The preliminary user test has yielded encouraging results. The participants reported that the fixation-based initiation of the assistant was intuitive and natural, appreciating the silent and hands-free nature of the activation. Feedback highlighted the importance of carefully tuning the dwell time threshold and the acknowledgment feedback design. These initial findings suggest that this approach holds potential for creating more seamless and less intrusive interactions with AI assistant in a smart AR eyewear that is equipped every day.
@@ 3. Our Approach: Eye Fixation As The Trigger of Virtual AI Assistant
Our methodology utilizes eye tracking to ascertain the user’s gaze orientation and detects the interaction between where the user intends to direct their attention and a virtual avatar situated within a mixed reality environment. This approach is motivated by the typical initiation process in natural human communication, where an individual shifts their gaze from their current task to directly look at another person, facilitating conversation initiation and reducing misunderstandings. The virtual avatar is placed above the user’s line of sight, nearly out of the visual field, with a small segment still visible to signal the position of the AI assistant (see Figure 1). To activate the AI assistant, embodied by the virtual avatar, the user needs to fixate on the avatar for approximately two seconds, this duration was established through an informal test. This requirement effectively minimizes accidental activations due to brief glances. Upon activation, the user can verbally interact with the avatar without the necessity of using wake words or a physical switching mechanism.
@@ 5. Preliminary User Feedback and Insights
These initial feedback might suggest that fixation-based initiation could be a less intrusive way to engage with AI assistants, particularly when AI-powered AR glasses become commonplace. Future research will aim to conduct in-depth user studies, explore diverse interaction modalities, and incorporate context awareness. We believe that this study has the potential to enhance human-AI interaction frameworks within the rapidly developing field of augmented reality, addressing the limitations of conventional voice and button interfaces by providing a quiet, hands-free, and more natural interaction method.

### v2

@@ Abstract.
Engaging with AI assistants to gather essential information in a timely manner is becoming increasingly common. Traditional activation methods, like wake words such as Hey Siri, Ok Google, and Hey Alexa, are constrained by technical challenges such as false activations, recognition errors, and discomfort in public settings. Similarly, activating AI systems via physical buttons imposes strict interaction limitations as it demands particular physical actions, which hinders fluid and spontaneous communication with AI. Our approach employs eye-tracking technology within AR glasses to discern a user’s intention to engage with the AI assistant. By sustaining eye contact on a virtual AI avatar for a specific time, users can initiate an interaction silently and without using their hands. Preliminary user feedback suggests that this technique is relatively intuitive, natural, and less obtrusive, highlighting its potential for integrating AI assistants fluidly into everyday interactions.
@@
journalyear: 2025copyright: rightsretainedconference: The Augmented Humans International Conference 2025; March 16–20, 2025; Masdar City, Abu Dhabi, United Arab Emiratesbooktitle: The Augmented Humans International Conference 2025 (AHs ’25), March 16–20, 2025, Masdar City, Abu Dhabi, United Arab Emiratesdoi: 10.1145/3745900.3746107isbn: 979-8-4007-1566-2/25/03ccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Interaction techniquesccs: Computing methodologies Natural language processingccs: Applied computing
@@ 1. Introduction
The preliminary user test has yielded encouraging results. The participants reported that the fixation-based initiation of the assistant was intuitive and natural, appreciating the silent and hands-free nature of the activation. Feedback highlighted the importance of carefully tuning the dwell time threshold and the acknowledgment feedback design. These initial findings suggest that this approach holds potential for creating more seamless and less intrusive interactions with AI assistant in smart AR eyewear that is equipped every day.
@@ 3. Our Approach: Eye Fixation As The Trigger of Virtual AI Assistant
Our methodology utilizes eye tracking to ascertain the user’s gaze orientation and detects the interaction between where the user intends to direct their attention and a virtual avatar situated within a mixed reality environment. This approach is motivated by the typical initiation process in natural human communication, where an individual shifts their gaze from their current task to directly look at another person, facilitating conversation initiation and reducing misunderstandings. The virtual avatar is placed above the user’s line of sight, nearly out of the visual field, with a small segment still visible to signal the position of the AI assistant (see Figure 1). To activate the AI assistant, embodied by the virtual avatar, the user needs to fixate on the avatar for approximately two seconds. This duration was established through an informal test. This requirement effectively minimizes accidental activations due to brief glances. Upon activation, the user can verbally interact with the avatar without the necessity of using wake words or a physical switching mechanism.
@@ 5. Preliminary User Feedback and Insights
This initial feedback might suggest that fixation-based initiation could be a less intrusive way to engage with AI assistants, particularly when AI-powered AR glasses become commonplace. Future research will aim to conduct in-depth user studies, explore diverse interaction modalities, and incorporate context awareness. We believe that this study has the potential to enhance human-AI interaction frameworks within the rapidly developing field of augmented reality, addressing the limitations of conventional voice and button interfaces by providing a quiet, hands-free, and more natural interaction method.